OpenStudy (anonymous):

determine the axis of symmetry of the quadratic equation y=4x^2+10x+5

OpenStudy (anonymous):

Remember, the axis of symmetry is the line that cuts down the quadratic. Correct?

OpenStudy (anonymous):

The line also goes through the vertex (a very interesting, yet logical fact). Remember the vertex is the maximum or minimum of a quadratic, the point at which the graph's direction changes

OpenStudy (anonymous):

The vertex of this quadratic is -b/2a, where b=10, and a = 4. -10/8 = -5/4. The axis of symmetry is the vertical line, thus it's x=-5/4